### v1

# Search for a gravitational wave background from primordial black hole binaries using data from the first three LIGO-Virgo-KAGRA observing runs

Abstract
Using the cross-correlation data from the first three observing runs of the LIGO-Virgo-KAGRA Collaboration, we search for a gravitational-wave background (GWB) from primordial black holes, arising from the superposition of compact binary coalescence events. We consider both early and late binary formation mechanisms, and perform Bayesian parameter inference, investigating different prior distributions of the model parameters. From the non-detection of the GWB, we provide constraints on the fraction of primordial black holes contributing to the present dark matter energy density.
preprint: IFT-UAM/CSIC-24-171, KCL-PH-TH-2024-64
Introduction.— Gravitational-wave (GW) astronomy has the potential to revolutionize our understanding of the Universe by providing insights into its most energetic and dynamic phenomena. GW detections allow us to explore regions of space-time that are inaccessible through electromagnetic observations. Since the first binary black hole (BBH) merger detected by the LIGO-Virgo-KAGRA (LVK) collaboration [1], the field has rapidly expanded with numerous GW events reported over the past decade [2, 3, 4]. These observations have advanced our knowledge of compact objects, their merger mechanisms, and have enabled tests of general relativity in the strong-field regime [5]. Additionally, GW signals have been used to measure the Hubble constant through standard sirens [6].
Beyond individual events, the gravitational wave background (GWB)—a collective signal from many unresolved sources—provides a powerful tool for probing both astrophysical and cosmological processes. Astrophysical contributions to the GWB arise from unresolved BBH, binary neutron star (BNS), and black hole-neutron star (BHNS) mergers throughout cosmic history [7]. These compact binaries may form via various channels, including isolated binary evolution and dynamical interactions in dense stellar environments. On the cosmological side, potential GWB sources include phase transitions in the early universe [8, 9], cosmic strings [10, 11], and primordial fluctuations from inflation [12]. Detecting a cosmological GWB would open a new window into the physics of the early universe and high-energy phenomena beyond the reach of particle accelerators [13, 14].
In this article, we focus on primordial black holes (PBHs), which are hypothesized to form from the collapse of large density fluctuations in the early universe [15]. Unlike astrophysical black holes, which are remnants of massive stars, PBHs can span a wide range of masses, from asteroid-like to supermassive scales, depending on the model. PBHs are gaining particular attention as they offer unique insights into conditions in the early universe, such as the spectrum of primordial density fluctuations, phase transitions, and inflationary physics. Moreover, PBHs are considered potential candidates for dark matter [16, 17, 18, 19] and for observed compact binary coalescences (CBCs) [20, 21, 22].
Several previous studies have investigated the possibility that some of the detected CBC events originate from PBH mergers [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 35, 41, 42, 43]. These studies typically analyze the observed population of CBCs obtained through individual detections by examining their mass distributions, spin orientations, and redshift distributions to discuss whether the properties of these events are consistent with an astrophysical origin or suggest a primordial origin. For example, a BBH merger with component masses in the so-called “mass gap” region—where standard stellar evolution models predict a scarcity of black holes due to pair-instability supernovae—has prompted speculation about a primordial origin [44, 32, 39, 32, 45, 31]. Furthermore, a subsolar mass event, can serve as distinct evidence of PBHs [46, 47, 48].
These analyses compare the observed GW event rates with the predicted merger rates from PBH populations. Note, however, that LVK binary abundance constraints rely on calculations showing that PBHs form binaries very efficiently in the early universe, treating PBHs as constant Schwarzschild masses. This assumption is not well justified, since during such early times the Hubble horizon is still small. Consequently, one cannot ignore the fact that these PBHs are embedded in an expanding background [49].
Another approach is to search unresolvable binaries as a GWB. Some studies have explored constraints on PBHs by considering the GWB produced by their mergers, a line of research pioneered in [50]. Constraints have been discussed in the literature [26, 28, 51, 52, 53, 54, 55], involving typically a rough comparison between the theoretical predictions and the upper bounds obtained from the GWB search performed by the LVK collaboration [56, 57, 58], which assumes a power-law spectrum $\Omega_{\rm GW}\propto f^{\alpha}$. However, for GWB generated by compact binaries with mass distributions extending beyond $\sim 50M_{\odot}$, the spectrum deviates from a simple power law.
Motivated by this, we analyze data from the first three LVK observing runs (O1–O3) [56, 57, 58] and perform Bayesian parameter inference for the GWB from PBHs with a wide range of masses, including sub-solar masses up to $10^{3}$ solar masses. This allows us to account for the complete spectral shape and explore a broad parameter space that extends beyond the detection of individual mergers by LVK. A recent work in this direction is [43], although it is primarily focused on distinguishing between astrophysical and primordial BBHs. In addition, the PBH model used in [43] was limited to a fixed mass of $1$ and $30M_{\odot}$. It is also important to note that, for the first time, we consider in our study the combination of early and late binary formation channels to get corresponding constraints.
In what follows, we aim to effectively bridge the gap between theoretical studies and observational data through robust Bayesian analysis. Using the most up-to-date publicly available data from the LVK (O1-O3) searches, we provide more reliable and comprehensive constraints on PBH abundance for peaked mass distributions across a broad mass range, a topic gaining significant attention in the field, particularly in relation to other cosmological constraints.
Modeling.— To parameterize the abundance of PBH, we commonly use $f_{\rm PBH}=\Omega_{\rm PBH}/\Omega_{\rm DM}$, which is the fraction of PBH in the dark-matter energy density today. One of the key factors influencing the GW spectrum in our model is the mass function of PBHs. It is commonly parametrized using monochromatic or Log-Normal distributions. They may not accommodate a mass function with a detailed shape, such as the one associated with the QCD phase transition [59, 60, 61], but it is a good approximation for PBHs arising from a peak in the primordial power spectrum [62].
Using the PBH energy density $\rho_{\rm PBH}$, we define the Log-Normal PBH mass function $p(m)$ as
$p(m)=\frac{1}{\rho_{\rm PBH}}\frac{\mathrm{d}\rho_{\rm PBH}}{\mathrm{d}\ln m}.$ (1)
For a Log-Normal Gaussian distribution of the PBH mass function, we have
$p(m)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left[-\frac{(\ln m-\ln\mu)^{2}}{2\sigma^{2}}\right].$ (2)
The mass function is normalized to satisfy $\int p(m)\mathrm{d}\ln m=1$.
The amplitude of the GWB is often characterized by the dimensionless parameter
$\Omega_{\rm GW}\equiv\frac{1}{\rho_{c,0}}\frac{\mathrm{d}\rho_{\rm GW}}{\mathrm{d}\ln f},$ (3)
where $\rho_{c,0}=3H_{0}^{2}/(8\pi G)$ is the critical density of the Universe today. The GWB formed by the ensemble of CBC events can be calculated by summing up the energy spectrum of each binary system, divided by $(1+z)$ to account for the redshifting of the gravitons since emission and taking into account the merger rate distribution [63]
$\Omega_{\rm GW}(f)$ $=\frac{f}{\rho_{c,0}}\int_{0}^{z_{\rm max}}\mathrm{d}z\int\mathrm{d}\ln m_{1}\ \mathrm{d}\ln m_{2}\ \frac{p(m_{1})p(m_{2})}{(1+z)H(z)}$
×$\times\frac{d^{2}R_{\rm EB/LB}}{d\ln m_{1}d\ln m_{2}}\frac{\mathrm{d}E_{\rm GW}}{\mathrm{d}f_{\rm r}},$ (4)
where $f_{\rm r}=(1+z)f$ is the GW frequency in the source frame. Here $\frac{d^{2}R_{\rm EB/LB}}{d\ln m_{1}d\ln m_{2}}$ describes the differential merger rate per unit time, comoving volume, and mass interval where EB and LB stand for “Early Binary” and “Late Binary” respectively and denote different binary formation mechanisms described in the next subsection. $\mathrm{d}E_{\rm GW}/\mathrm{d}f_{r}$ is the single source energy spectrum, which is a function of the masses of the compact objects, $m_{1}$ and $m_{2}$; its fitting function including inspiral, ringdown, and merger phases is given in [64].
Early binary formation.— Early binaries are formed in the radiation-dominated era shortly after PBH formation. In this scenario, a binary is formed from a pair of closely spaced PBHs with the surrounding third object acting on the pair via a tidal force that generates the angular momentum of the binary [65, 21]. The merging rate (per unit logarithmic mass of the two binary components) of early binaries at a time $t$ is given by [66, 27, 24, 21, 67].
$\frac{d^{2}R_{\rm EB}}{d\ln m_{1}d\ln m_{2}}$ $=\frac{1.6\times 10^{6}}{\rm Gpc^{3}yr}f_{\rm PBH}^{53/37}\Big{[}\frac{t}{t_{0}}\Big{]}^{-34/37}$
×$\times\left(\frac{m_{1}+m_{2}}{M_{\odot}}\right)^{-32/37}\left[\frac{m_{1}m_{2}}{(m_{1}+m_{2})^{2}}\right]^{-34/37}$
×$\times S(m_{1},m_{2},f_{\rm PBH}),$ (5)
where $m_{1}$ and $m_{2}$ are the two black hole masses and $t_{0}$ is the age of the Universe. A rate suppression factor, $S$, has been introduced to account for two principal mechanisms that impede binary formation due to local matter inhomogeneities and nearby PBHs $S_{1}(m_{1},m_{2},f_{\rm PBH})$ and the clustering due to their initial Poissonian fluctuations $S_{2}(f_{\rm PBH})$ [27, 28].
Analytical prescriptions for these suppression factors have been studied in [28, 27]. We present these expressions in the Appendix A, including the ranges relevant to our work. Notably, for narrow mass distributions with $\sigma<1$, the suppression factor $S$ is less than 0.2 (see Fig. 2). In this study, we fix $S$ as $S=S_{1}\times S_{2}=0.1$. This choice is reasonable in the range $0.010.001<f_{\rm PBH}<$ 0.01, where our constraints are most applicable. We have confirmed that our specific choice of $S$ does not significantly affect our results. As detailed in the Appendix A, the suppression factor has significant uncertainties, particularly in the case of broad mass distributions, which led us to focus on the scenario of a narrow mass distribution in this paper.
Late binary formation.— Late binaries form much later due to dynamical capture in the clusters of PBHs during the matter-dominated era [20]. For a general mass distribution of PBHs, an expression for the merger rate has been proposed in the literature [68, 69]:
$\frac{d^{2}R_{\rm LB}}{d\ln m_{1}d\ln m_{2}}=\frac{R_{\rm clust}}{\rm Gpc^{3}yr}f_{\rm PBH}^{2}\frac{(m_{1}+m_{2})^{10/7}}{(m_{1}m_{2})^{5/7}}\,,$ (6)
where $R_{\rm clust}=400$ (see the discussion in [31] for this choice) is a scaling factor that encapsulates the clustering properties and velocity distribution of the PBHs. This relationship is derived from the two-body capture process occurring within a cluster and operates under the assumption that the time required for the binary system to merge is significantly shorter than the age of the universe.
Large uncertainties remain on the exact value of $R_{\rm clust}$ and its possible mass and PBH model dependence. In our case, the rate of late binaries remains subdominant compared to early binaries, except for $m_{\rm PBH}\gtrsim 200M_{\odot}$. Using GWB searches to set new limits on the PBH abundance for a large mass range is an interesting avenue that one could follow using the methodology we present here. One can note that the limits obtained on $f_{\rm PBH}$ using our benchmark value of $R_{\rm clust}$ can be redrawn for any other value, with the simple rescaling $f_{\rm PBH}\longleftrightarrow f_{\rm PBH}(400/R_{\rm clust})^{1/2}$. This way, the limits we obtain can in fact become much more stringent in the case of smaller and denser PBH clusters.
Methodology.— Let us outline the Bayesian framework employed in our search for GWB originating from PBHs and other compact binary sources. Our methodology was first introduced in [70] and closely follows that of previous studies [71, 72, 73, 74, 75, 76, 77, 78, 79, 57, 58], with specific adaptations to our analysis.
We utilize data from the Advanced LIGO detectors located in Hanford (H) and Livingston (L), as well as data from the Virgo (V) observatory, during the first three observing runs (O1–O3) [80, 81, 82, 83]. Our search incorporates all available baselines (HL, HV, LV) (for O1 and O2 we use only HL) to maximize sensitivity and leverage the complementary geographical locations of the detectors.
Following the notation of [75], the primary observable is the cross-correlation estimator $\hat{C}_{IJ,f_{i}}$ which measures the correlation between detectors $I$ and $J$ as a function of discrete frequency $f_{i}$. This estimator is constructed by cross-correlating the discrete Fourier transforms of the outputs from different detectors. It is then transformed into the estimator $\hat{\Omega}^{IJ}_{{\rm GW},f_{i}}$ by multiplying with the relevant numerical factors and the overlap reduction function. These definitions are given in Appendix B. The likelihood function for the search of an isotropic GWB is [70, 75]
$p(\hat{\Omega}^{IJ}_{{\rm GW},f_{i}}|\boldsymbol{\theta},\lambda)\propto\exp\left[-\frac{1}{2}\sum_{IJ}\sum_{f_{i}}\left(\frac{\hat{\Omega}^{IJ}_{{\rm GW},f_{i}}-\lambda\Omega_{\rm GW}(f_{i}|\boldsymbol{\theta})}{\sigma^{IJ}_{f_{i}}}\right)^{2}\right]\,,$ (7)
where $\sigma^{IJ}_{f_{i}}$ denotes the variance of the cross-correlation estimator, which is derived from the noise power spectral density of each detector and $\Omega_{\rm GW}(f_{i}|\boldsymbol{\theta})$ represents the theoretical model to be compared with the data, characterized by a set of parameters $\boldsymbol{\theta}$. The sum runs over the frequency bins $f_{i}$ as well as different pairs of the detectors labeled by $I,J={\rm H,L,V}$. We then construct the posterior distribution using Bayes’ Theorem,
$p(\boldsymbol{\theta},\lambda|\hat{\Omega}^{IJ}_{{\rm GW},f_{i}})=\frac{p(\hat{\Omega}^{IJ}_{{\rm GW},f_{i}}|\boldsymbol{\theta},\lambda)p(\theta,\lambda)}{p(\hat{\Omega}^{IJ}_{{\rm GW},f_{i}})}\,.$ (8)
[…]
Prior Distributions.— We employ two distinct sets of priors in our analysis. The first set, referred to as Agnostic Log-Normal, includes the parameters: the mean of the Log-Normal mass distribution $\mu$, its width $\sigma$, and the PBH fraction $f_{\rm PBH}$ in the dark matter energy density today, as introduced previously. The second set of priors, termed Mono-chromatic (MN), is identical to the Agnostic Log-Normal (ALN) priors except that the width $\sigma$ is fixed at $10^{-2}\,M_{\odot}$.
Figure 1: Two-sigma ($2\sigma$) constraints on the PBH fraction ($f_{\rm PBH}$) as a function of $M_{\rm PBH}$($=\mu$) from the search for the gravitational-wave background (GWB) using the ALN (MN) constraint is shown in solid (dashed) red. We also display supernova lensing constraints (SNe) [86] in purple, Massive Compact Halo Object (MACHO) constraints [87] in orange designated as (EROS), the recent Optical Gravitational Lensing Experiment (OGLE) constraints are shown in green with different Milky Way (MW) models [88], and cosmic microwave background (CMB) constraints from [89]. We used the PBHBounds [90] to generate this plot.
For both sets we also take into account for an additional CBC background from astrophysical black holes. Given their typical mass range of ${\cal O}(1-10)M_{\odot}$, the spectrum is well approximated by a simple power law characterized by a single parameter, and is expressed as $\Omega_{\rm cbc}$ as $\Omega_{\rm CBC}(f)=\Omega_{\rm cbc}(f/f_{\rm ref})^{2/3}$, with $f_{\rm ref}=25$ Hz [91]. Therefore, our model is $\Omega_{\rm GW}=\Omega_{\rm CBC}+\Omega_{\rm EB}+\Omega_{\rm LB}$.
The prior distributions for both sets of model parameters are specified in Table  1. These ranges encompass plausible values based on theoretical models and existing constraints [26, 91, 28]. Note that we focus on relatively sharp mass distributions due to uncertainties in the theoretical predictions for broad mass distributions as already mentioned. Thus, the maximum value of $\sigma$ for ALN prior is set to 1.
Bayesian Search.— In order to perform Bayesian parameter inference, we employ the nested sampling algorithm from the dynesty package [92] for robust parameter space exploration. To expedite the computation of integrals in the gravitational wave spectrum calculations, we utilize Sobol Monte Carlo integration [93], a quasi-random low-discrepancy sequence method that improves convergence rates over traditional Monte Carlo approaches, thereby reducing computational overhead. To account for calibration uncertainties in GW detector data, we use the Python library pygwb [75], which integrates these uncertainties into the parameter estimation process.
Parameter ALN Prior MN Prior
$\Omega_{\rm cbc}$ $\mathcal{LU}$($10^{-10}$, $10^{-7}$) $\mathcal{LU}$($10^{-10}$, $10^{-7}$)
$\mu\,(M_{\odot})$ $\mathcal{LU}$($10^{-1}$, $10^{3}$) $\mathcal{LU}$($10^{-1}$, $10^{3}$)
$\sigma$ $\mathcal{LU}$($10^{-2}$, $10^{0}$) Fixed at $10^{-2}$
$f_{\rm PBH}$ $\mathcal{LU}$($10^{-5}$, $1$) $\mathcal{LU}$($10^{-5}$, $1$)
Table 1: Prior distributions for the model parameters in the ALN and MN frameworks used in the Bayesian analysis. $\mathcal{LU}$ denotes a LogUniform distribution. For the MN framework, $\sigma$ is fixed at $10^{-2}$.
Our analysis is based on the combined cross-correlation dataset from the first three observing runs of the LVK collaboration. We applied data quality cuts and vetoes following the procedures outlined in previous isotropic GWB searches by the LVK [56, 57, 58]. For parameter sampling, we utilized $5\times 10^{4}$ live points with a convergence criterion of $d\ln(z)<0.1$.
The results indicate that no significant GWB associated with PBHs or CBCs has been detected over the noise hyphothesis, with logarithmic Bayes factors $\sim-1$. We establish upper limits on the CBC energy density parameter, $\Omega_{\rm cbc}$, of $6.0\times 10^{-9}$ for ALN priors and $6.2\times 10^{-9}$ for MN priors at the $2\sigma$ level which are consistent with LVK contraints [58].
From the non-detection, we can constrain the fraction of PBHs, $f_{\rm PBH}$, as a function of the mean mass $\mu$. The obtained constraint, along with other studies from lensing surveys like EROS ([87]), OGLE ([88]), supernova Ia ([86]) and cosmic microwave background (CMB) constraints are summarized in Fig. 1. We also provide the resulting corner plots for all model parameters in Appendix C.
The constraint on $f_{\rm PBH}$ is obtained by marginalizing over the mass function width $\sigma$ in the ALN scenario. However, because of the relatively narrow range of $\sigma$, our analysis shows no significant differences between the two distinct prior sets when imposing constraints on $f_{\rm PBH}$.
These findings demonstrate that GWB constraints are and will continue to be a vital probe, particularly within the mass range $[10^{1},3\times 10^{2}]\,M_{\odot}$. This particular mass range is dominated by the early binaries and our constraints remain between $0.010.001<f_{\rm PBH}<$ 0.01. For $M_{\rm PBH}>10^{2}M_{\odot}$, the GWB spectral amplitude is dominated by the late binary formation channel. Although the LVK loses its constraining power sharply in this mass range due to the limited frequency range, this suggests an interesting possibility for constraining this mass range using the late binary formation channel with the Einstein Telescope [94, 95] and Cosmic Explorer[96, 97], which have greater sensitivity at low frequencies.
Conclusion.— In this study, we conducted a targeted search for a GWB originating from PBHs using cross-correlation data from the first three observing runs of the LVK Collaboration [56, 57, 58]. Unlike previous research that focused on the population of resolved gravitational-wave events, we utilized cross-correlated LVK data specifically to search for a GWB resulting from the superposition of many unresolved binary events. Our mass range encompasses PBH binaries with component masses both below and above those detected in the LVK catalog [2, 4].
Our approach advances beyond previous stochastic analyses that utilized the LVK upper limit derived under the assumption of a power-law shape for the spectrum. By directly modeling the expected GWB signal from PBH mergers, we present, for the first time, a direct GWB search specifically targeting PBH mergers.
Although we did not find significant evidence for a GWB from PBHs, we established new constraints on the fraction of PBHs contributing to the current dark matter energy density, $f_{\rm PBH}$, as a function of PBH mass assuming a log-normal mass distribution. These constraints complement existing bounds from individual CBC events [26, 28, 35] and other observational methods such as microlensing and cosmic microwave background observations [86, 87, 88, 89].
Future constraints will undoubtedly provide deeper and broader insights into the abundance of PBHs across a wider mass range and will offer valuable information for pinpointing the origin of black holes. Additionally, the broad mass function, in particular those predicted by QCD phase transitions, remains an interesting case for investigation. In future work, we plan to apply our analysis to this model using more accurate descriptions of their merger rates that are currently under development.
Acknowledgments.— The authors would like to thank Vuk Mandic and the LVK Stochastic group for comments.
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The software packages used were matplotlib [98], numpy [99], scipy [100], dynesty [92], pygwb [75] and PBHBounds [90]. This manuscript was assigned LIGO-Document number LIGO-P2400579.
SC acknowledges support from the Belgian Fund for Research FNRS-F.R.S. through an Incentive Grant For Scientific Research and through the IISN convention 4.4501.19.
SK is supported by the Spanish Attraccion de Talento contract no. 2019-T1/TIC-13177 granted by the Comunidad de Madrid, the I+D grant PID2020-118159GA-C42 and the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S funded by MCIN/AEI/10.13039/501100011033, the Consolidación Investigadora 2022 grant CNS2022-135211, and Japan Society for the Promotion of Science (JSPS) KAKENHI Grant no. 20H01899, 20H05853, and 23H00110.
MS acknowledges support from the Science and Technology Facility Council (STFC), UK, under the research grant ST/X000753/1.
Appendix C: Corner Plots.—
Figure 3: Corner plots for the ALN (top) MN (bottom) posteriors. Darker regions are at $1\sigma$ and lighter regions are at $2\sigma$ level.